OpenStudy (anonymous):

true and false! If the dimensions of a circle are tripled, the area of the resulting circle will also be three times larger True or False If a circle is enlarged to have a circumference that is four times the original, then the original radius was multiplied by four. true or false If a parallelogram is reduced to have dimensions that are half as long, then the perimeter is one fourth the original length True or false?

jimthompson5910 (jim_thompson5910):

were you able to get anywhere?

OpenStudy (anonymous):

No

jimthompson5910 (jim_thompson5910):

A = pi*r^2 ... original area A = pi*(3r)^2 ... triple the radius A = pi*9r^2 A = 9(pi*r^2) so the area is not tripled, but multiplied by 9

jimthompson5910 (jim_thompson5910):

so saying "If the dimensions of a circle are tripled, the area of the resulting circle will also be three times larger" is false the area isn't tripled

OpenStudy (anonymous):

awe thank you!

jimthompson5910 (jim_thompson5910):

saying "If a circle is enlarged to have a circumference that is four times the original, then the original radius was multiplied by four." is true because C = 2*pi*r ... original circumference 4*C = 4*2*pi*r ... quadruple the circumference (ie multiply both sides by 4) 4*C = 2*pi*(4r) ... and we can see that the radius has been multiplied by 4

jimthompson5910 (jim_thompson5910):

I'll let you take on the last one. Tell me what you get and why you think that is the answer

OpenStudy (anonymous):

The last one is false